OpenStudy (anonymous):

Thanks for helping! Help with symmetries! Identify, graph, and state the symmetries of each polar equation. Write the scale that you are using for the polar axis. r=9cos(5 theta)

OpenStudy (anonymous):

@Hero @jim_thompson5910 @phi @zepdrix @SithsAndGiggles @wio @sleepyjess @Data_LG2

OpenStudy (mathmate):

Hint: To make an x-y plot, use \(x=r~\cos(\theta)\) \(y=r~\sin(\theta)\)